# G2D: Boosting Multimodal Learning with Gradient-Guided Distillation

### v1

@@ 4 Experiments
We evaluate the G2D framework on the basis of the following questions: Q1: How does G2D compare to state-of-the-art methods for addressing modality imbalance and overall multimodal performance in supervised tasks? Q2: How does G2D influence the modality gap in multimodal learning?, Q3: Does G2D enhance feature space alignment of unimodal and multimodal models?, Q4: Which fusion and suppression techniques are the best for G2D?, and Q5: How SMP is influencing the multimodal learning?
@@ 4.1.1 Experimental Setup
Datasets. We choose four multimodal classification datasets and one regression dataset. CREMA-D [3] is an audio-visual dataset for speech emotion recognition with six emotion classes. AV-MNIST [34] is a synthetic dataset with PCA-projected MNIST images and audio spectrograms for ten-digit classes. VGGSound [5] is a large-scale audio-visual dataset with 309 classes of everyday audio events, featuring video clips of 10 seconds each. UR-Funny [13] is a binary classification dataset for humor detection that incorporates text, visual gestures, and acoustic modalities. MIS-ME [28] dataset contains raw soil patch images and corresponding meteorological data for the soil moisture estimation task. To the best of our knowledge, we are the first to evaluate the modality imbalance in a multimodal regression setting and use tabular data for multimodal learning problems.
@@
Backbone and Hyperparameter Settings. For audio-visual datasets (CREMA-D, VGGSound, and AV-MNIST), we use ResNet-18 [14] as the encoder for both audio and video modalities in the teacher and student models. For the UR-Funny dataset, which involves audio, visual, and text modalities, we use a Transformer-based encoder [33] for each modality. For MIS-ME, we adopt MobileNetV2 [29] as the image feature extractor and use the fully connected neural network proposed in [28] for processing tabular meteorological data. To ensure a fair comparison, we use identical backbone architectures across all baseline models and employ late fusion for training. All models are optimized using SGD with a momentum of 0.9 and a weight decay of 0.0001. We use a batch size of 16 across all experiments and train on a single NVIDIA A10 GPU. More details on experimental settings are provided in Sec. A.4.
@@ 4.3 Ablation Study
Experimental setups of the ablation studies and ablation tables with unimodal performances can be found in Sec. A.5 and Sec. A.6, respectively.
@@
G2D with Various Fusion Modules. Table 5 compares the effect of different fusion strategies on G2D. Out of the traditional fusion methods used for the modality imbalance problem, late fusion consistently achieves the best results. This highlights the effectiveness of leveraging independent unimodal representations while preserving their distinct contributions. However, Cross-Attention based fusion closely follows the performance of late fusion, demonstrating its ability to enhance cross-modal interactions by dynamically attending to relevant features. Concat fusion provides strong performance but falls slightly behind Late Fusion. FiLM and BiGated offer adaptive feature integration yet fail to match the top-performing methods.
@@
Table 7: Effect of $\alpha$ and $\beta$ in G2D with SMP
Dataset ($\mathbf{\alpha}$, $\mathbf{\beta}$) weights
(0, 0) (0.25, 0.75) (0.5, 0.5) (0.75, 0.25) (1, 0) (0, 1) (1, 1)
CREMA-D 80.78 84.41 84.95 84.81 82.39 84.68 85.89
UR-Funny 63.58 64.79 64.29 64.29 64.59 64.69 65.49
Effect of $\alpha$ and $\beta$ in G2D. $\alpha$ and $\beta$ denote the weighting coefficients of feature loss and logit loss, respectively in the proposed G2D loss. Table 7 evaluates the effect of changing the weightage of feature loss and logit loss on G2D. Assigning full weight to both losses ($\alpha=1,\beta=1$) yields the best overall performance, highlighting their combined importance in multimodal learning. In contrast, removing both losses ($\alpha=0,\beta=0$) significantly reduces performance, confirming their necessity. While different weight combinations impact results, incorporating both losses with higher weight leads to greater improvements across datasets.
@@ A.2.7 MMPareto: Boosting Multimodal Learning with Innocent Unimodal Assistance [37]
A.2.8 On Uni-Modal Feature Learning in Supervised Multi-Modal Learning [7]
@@ A.2.9 ReconBoost: Boosting Can Achieve Modality Reconcilement [16]
A.2.10 Facilitating Multimodal Classification via Dynamically Learning Modality Gap [42]
Dynamically Learning Modality Gap (DLMG) addresses the modality imbalance problem in multimodal learning by focusing on disparities in category label fitting across different modalities. Unlike prior methods that primarily regulate learning rates or gradient contributions, DLMG leverages contrastive learning to align modality representations and reduce dominance effects. The approach dynamically integrates supervised classification loss and contrastive modality matching loss through either a heuristic strategy or a learning-based optimization strategy that adjusts their relative importance during training. By progressively refining modality alignment while maintaining label supervision, DLMG minimizes performance gaps between dominant and non-dominant modalities, leading to a more balanced and effective multimodal learning process.
@@ A.4.1 Model Architectures
Transformers are powerful architectures designed for handling sequential data and capturing long-range dependencies through self-attention mechanisms. In our implementation, we employ Transformers as encoders for the audio, video, and text modalities of the UR-Funny dataset. Following the approach described in [21], we utilized a 4-layer Transformer encoder with eight attention heads and a hidden dimension of 768 for each modality in the UR-Funny dataset. Input features were projected to a 768-dimensional embedding using a convolutional layer, ensuring consistency across different modalities. We employed a similar initialization strategy to ResNet-18 to facilitate stable training.
@@ A.4.2 Hyperparameters
A.5 Ablation Experimental Setup
We give four ablation studies in the main paper for the proposed G2D: (1) Analyze the impact of SMP in improving multimodal performance; (2) Evaluate G2D across multiple fusion modules; (3) Compare the effectiveness of partial vs complete suppression using SMP; (4) Analyze the impact of feature loss ($\alpha$) and distill loss ($\beta$). We follow the existing approach for conventional methods[24] and deactivate one modality for Late Fusion, FiLM, and BiGated, we deactivate one modality within the multimodal model.
To analyze fusion modules, we employ conventional methods (Sum, Concat, Late Fusion), modulation-based fusion techniques (FiLM, BiGated), and Cross-Attention. To measure unimodal performance in a multimodal setting, we follow [46]. Sum method involved feeding features from the evaluated modality into the corresponding network head prior to summation, and the Concat method used partitioned network heads to measure unimodal performance. For Late Fusion, FiLM, BiGated, and Cross-Attention we deactivate one modality within the multimodal model.
For the modality suppression study, we compared two strategies: partial suppression and complete suppression. Partial suppression was performed using the technique introduced in OGM-GE, which involved gradient modulation with $1-\tanh(x)$, where $x$ is the ratio of scores between two modalities. Complete suppression, on the other hand, applies sequential modality prioritization, allowing each suppressed modality to train to convergence while zeroing out gradients for all other modalities.
A.6 Ablation Tables of Main Paper with Unimodal Performance
Table 8: Performance comparison of G2D with SMP vs. without SMP on three audio-visual datasets reported in accuracy (%).
Type Without SMP With SMP
CREMA-D Audio 51.61 56.45
Visual 56.59 72.72
Multi 78.63 85.89
AV-MNIST Audio 39.62 39.10
Visual 65.18 65.09
Multi 72.76 73.03
VGGSound Audio 42.01 39.43
Visual 28.28 29.88
Multi 52.17 53.82
Table 9: Performance comparison of G2D with SMP vs. without SMP on the UR-Funny dataset across audio-video (A-V), audio-text (A-TXT), visual-text (V-TXT), and audio-visual-text (A-V-TXT) combinations, reported in accuracy (%).
Type Without SMP With SMP
A-V Audio 56.84 59.05
Visual 55.84 58.05
Multi 62.07 62.98
A-TXT Audio 55.73 59.86
Text 59.76 58.85
Multi 63.18 63.28
V-TXT Visual 53.92 56.34
Text 55.53 56.04
Multi 63.38 63.48
A-V-TXT Audio 52.92 59.15
Visual 52.62 55.94
Text 57.14 58.15
Multi 63.78 65.49
Dataset Sum Concat
Late
Fusion
FiLM BiGated
CREMA-D Audio 56.18 51.21 56.45 48.25 48.12
Visual 66.26 71.37 72.72 60.75 63.04
Multi 81.59 83.60 85.89 84.27 81.32
AV-MNIST Audio 39.87 35.73 39.10 41.44 42.40
Video 65.33 64.00 65.09 61.59 65.48
Multi 72.70 72.98 73.03 72.73 72.89
VGGSound Audio 34.96 38.87 39.43 19.09 30.35
Visual 27.28 30.35 29.88 23.12 28.53
Multi 50.67 53.40 53.82 48.11 46.66
UR-Funny Audio 50.70 55.13 59.15 51.41 44.27
Visual 49.30 55.63 55.94 44.27 49.30
Text 58.55 58.15 58.15 52.62 52.21
 Multi 63.08 64.49 65.49 63.48 63.38
Table 11: Partial vs. Complete modality suppression in G2D
Type Partial Complete
CREMA-D Audio 58.60 56.45
Visual 70.30 72.72
Multi 81.99 85.89
AV-MNIST Audio 35.70 39.10
Visual 57.71 65.09
Multi 72.83 73.03
VGGSound Audio 36.54 39.43
Visual 24.76 29.88
Multi 51.16 53.82
UR-Funny Audio 55.03 59.15
Visual 54.93 55.94
Text 58.25 58.15
Multi 63.68 65.49

### v2

@@ 4 Experiments
We evaluate the G2D framework on the basis of the following questions: Q1: How does G2D compare to state-of-the-art methods for addressing modality imbalance and overall multimodal performance in supervised tasks? Q2: How does G2D influence the modality gap in multimodal learning?, Q3: Does G2D enhance feature space alignment between unimodal and multimodal models?, Q4: How SMP is influencing the multimodal learning?, and Q5: Which fusion and suppression techniques are the best for G2D?
@@ 4.1.1 Experimental Setup
Datasets. We chose five multimodal classification datasets and one regression dataset. CREMA-D [4] is an audio-visual dataset for speech emotion recognition with six emotion classes. AV-MNIST [37] is a synthetic dataset with PCA-projected MNIST images and audio spectrograms for ten-digit classes. VGGSound [6] is a large-scale audio-visual dataset with 309 classes of everyday audio events, featuring video clips of 10 seconds each. UR-Funny [15] is a binary classification dataset for humor detection that incorporates text, visual gestures, and acoustic modalities. IEMOCAP [3] is an audio-visual-text dataset for emotion recognition in dyadic conversations. MIS-ME [30] is a regression dataset containing raw soil patch images and corresponding meteorological data for soil moisture estimation. To the best of our knowledge, we are the first to evaluate the modality imbalance in a multimodal regression setting and use tabular data for multimodal learning problems.
@@
Backbone and Hyperparameter Settings. For audio-visual datasets (CREMA-D, VGGSound, and AV-MNIST), we use ResNet-18 [16] as the encoder for both audio and video modalities in the teacher and student models. For the UR-Funny and IEMOCAP dataset, which involves audio, visual, and text modalities, we use a Transformer-based encoder [36] for each modality. For MIS-ME, we adopt MobileNetV2 [31] as the image feature extractor and use the fully connected neural network proposed in [30] for processing tabular meteorological data. To ensure a fair comparison, we use identical backbone architectures across all baseline models and employ late fusion for training. All models are optimized using SGD with a batch size of 16 and trained on a single NVIDIA A10 GPU. More details on experimental settings are provided in Sec. A.4.
@@ 4.2 Analysis of G2D
Further analysis of G2D, including distinction from key baselines and computational cost, is in Sec. A.6.
@@ 4.3 Ablation Study
G2D with Various Fusion Modules. Table 5 compares the effect of different fusion strategies on G2D. Out of the traditional fusion methods used for the modality imbalance problem, late fusion consistently achieves the best results. This highlights the effectiveness of leveraging independent unimodal representations while preserving their distinct contributions. However, Cross-Attention based fusion closely follows the performance of late fusion, demonstrating its ability to enhance cross-modal interactions by dynamically attending to relevant features. Concat fusion provides strong performance but falls slightly behind Late Fusion. FiLM and BiGated offer adaptive feature integration yet fail to match the top-performing methods. Descriptions of these fusion strategies are in Sec. A.3.
@@
Table 7: Effect of $\tau_{j}$ on G2D with two & three modalities
$(\tau_{1},\tau_{2})$ (0,150) (50,150) (100,150) (150,150)
CREMA-D 78.63 82.80 83.74 85.89
$(\tau_{1},\tau_{2},\tau_{3})$ (0,0,150) (50,50,150) (75,75,150)
IEMOCAP 75.30 76.99 77.19
Effect of Prioritization Epochs ($\boldsymbol{\tau_{j}}$). We analyze the impact of $\tau_{j}$, the hyperparameter controlling the epochs for each SMP stage, in Table 7. For instance, with two modalities ($k=2$), the schedule $(\tau_{1},\tau_{2})$ denotes epochs for training the weakest modality alone, followed by a joint training phase for both. Similarly, for three modalities ($k=3$), $(\tau_{1},\tau_{2},\tau_{3})$ defines epochs for the weakest alone, then the second weakest alone, and finally a joint phase for all three. The results in Table 7 show that systematically increasing the dedicated epochs for weaker modalities consistently improves G2D’s performance. This finding strongly validates our Hypothesis 1 (Sec. 3.3) that providing weaker modalities with dedicated, interference-free training phases is crucial for mitigating modality imbalance.
Additional ablations are provided in Sec. A.7.
@@ A.1.4 UR-Funny [15]
A.1.5 IEMOCAP [3]
The Interactive Emotional Dyadic Motion Capture (IEMOCAP) dataset is a widely used resource for emotion recognition, containing approximately 12 hours of audio-visual data from ten actors engaged in scripted and improvised dyadic conversations to elicit a range of authentic emotions. The dataset is rich in modalities, providing synchronized video, speech, facial motion capture data, and text transcriptions. Each utterance is annotated by multiple raters for both categorical emotions and dimensional attributes (valence, arousal, and dominance), making it suitable for diverse and nuanced modeling tasks. Due to its naturalistic dyadic interactions, IEMOCAP serves as a benchmark for developing emotion-aware conversational AI and understanding complex multimodal emotional cues.
@@ A.2.7 MMPareto: Boosting Multimodal Learning with Innocent Unimodal Assistance [40]
A.2.8 On Uni-Modal Feature Learning in Supervised Multi-Modal Learning (UMT) [8]
@@ A.2.9 ReconBoost: Boosting Can Achieve Modality Reconcilement [18]
A.2.10 Facilitating Multimodal Classification via Dynamically Learning Modality Gap (DLMG) [45]
DLMG addresses the modality imbalance problem in multimodal learning by focusing on disparities in category label fitting across different modalities. Unlike prior methods that primarily regulate learning rates or gradient contributions, DLMG leverages contrastive learning to align modality representations and reduce dominance effects. The approach dynamically integrates supervised classification loss and contrastive modality matching loss through either a heuristic strategy or a learning-based optimization strategy that adjusts their relative importance during training. By progressively refining modality alignment while maintaining label supervision, DLMG minimizes performance gaps between dominant and non-dominant modalities, leading to a more balanced and effective multimodal learning process.
A.2.11 Detached and Interactive Multimodal Learning (DI-MML) [10]
DI-MML proposes that modality competition is a direct result of the uniform learning objective used in traditional joint training frameworks. To eliminate this competition, DI-MML proposes a detached learning framework where each modality’s encoder is trained separately with its own isolated learning objective. To enable cross-modal interaction without reintroducing competition, the framework employs two key strategies: (1) a shared classifier is used to align features from different modalities into a common embedding space, and (2) a novel Dimension-decoupled Unidirectional Contrastive (DUC) loss is introduced. The DUC loss identifies “effective” and “ineffective” feature dimensions within each modality and then transfers knowledge unidirectionally from the effective dimensions of one modality to the corresponding ineffective dimensions of another. This strategy facilitates the exchange of complementary information while preserving the integrity of each modality’s well-learned features.
@@ A.4.1 Model Architectures
Transformers are powerful architectures designed for handling sequential data and capturing long-range dependencies through self-attention mechanisms. In our implementation, we employ Transformers as encoders for the audio, video, and text modalities of the UR-Funny and IEMOCAP dataset. Following the approach described in [23], we utilized a 4-layer Transformer encoder with eight attention heads and a hidden dimension of 768 for each modality in the UR-Funny and IEMOCAP dataset. Input features were projected to a 768-dimensional embedding using a convolutional layer, ensuring consistency across different modalities. We employed a similar initialization strategy to ResNet-18 to facilitate stable training.
@@ A.4.2 Hyperparameters
Table 8: Comparing G2D with DI-MML
Method Joint-Train  DI-MML G2D
CREMA-D 67.47 83.51 85.89
AV-MNIST 69.77 71.35 73.03
A.5 Comparison of G2D with DI-MML
In response to reviewer feedback, we provide an additional comparison against the state-of-the-art baseline DI-MML [10]. As shown in Table 8, G2D outperforms DI-MML on both the CREMA-D and AV-MNIST datasets. This result further validates the effectiveness of G2D in relation to current leading methods in the field.
A.6 Further Analysis of G2D
A.6.1 Distinction of G2D from UMT and OGM-GE
The primary novelty of G2D arises from its unique G2D loss and its Sequential Modality Prioritization (SMP) technique, and critically, from their synergistic combination.
Our G2D loss improves upon the distillation strategy of UMT [8] by incorporating a KL divergence-based logit loss. This addition is crucial for enabling the student model to learn the nuanced inter- and intra-class relationships captured by the unimodal teachers’ soft logits.
Furthermore, our SMP technique is fundamentally different from the gradient modulation in OGM-GE [26] in two principal ways:
  1. 1.
    Guidance for Modulation: OGM-GE calculates modality confidence from the student’s own encoders during training. This signal can be noisy and unreliable, especially in early stages. In contrast, SMP leverages stable confidence scores from pre-trained unimodal teachers, providing a more robust and accurate signal to identify weaker modalities automatically.
  2. 2.
    Suppression Mechanism: OGM-GE uses functions like $1-\tanh(\cdot)$ to only partially suppress dominant modalities, meaning they continue to train simultaneously and modality competition can persist. SMP enforces a complete gradient shutdown for non-prioritized modalities. This ensures that the prioritized weak modality trains in true isolation, more effectively mitigating interference from dominant modalities.
The results in Table 9 validate these distinctions, showing that the G2D loss alone surpasses UMT, SMP alone surpasses OGM-GE, and their combination yields the best overall performance.
Table 9: Comparing Components of G2D with UMT & OGM-GE
Method
Joint-
Train
UMT
G2D
Loss
OGM-
GE
SMP
G2D
(SMP + G2D Loss )
CREMA-D 67.47 67.61 78.63 72.18 80.78 85.89
AV-MNIST 69.77 72.33 72.76 71.08 72.51 73.03
A.6.2 Synergy of Distillation and Sequential Modality Prioritization
The motivation for integrating our G2D loss (via KD) with SMP is to address the limitations of using either technique alone. The distillation component leverages unimodal teachers—trained in isolation—to provide the student with stable, competition-free feature and logit targets. This guides the student towards more balanced representations than learning solely from GT labels amidst modality competition.
However, even with this guidance, the student’s modality encoders are still optimized simultaneously, which allows modality imbalance to persist. SMP is introduced to solve this. The crucial synergy lies in the fact that during the isolated training phases enforced by SMP, the prioritized weak modality learns not only from the ground-truth labels but also from the rich, interference-free knowledge distilled from its unimodal teacher via the G2D loss. This focused, dual-signal learning in an isolated context enables the robust development of weaker modalities. By combining SMP with our distillation objective, G2D mitigates modality imbalance more thoroughly and effectively than using either technique independently.
Table 10: Single-Batch Resource Metrics on CREMAD
Method     Total Memory (MB)     Total Execution Time (ms)
Joint-Train     12.1366 MB     4531.8
G2D     12.1998 MB     4539.8
A.6.3 Computational Cost
The training overhead of G2D is negligible, and there is no additional overhead during inference. This efficiency stems from the framework’s design: unimodal teacher models are pre-trained, and their outputs (e.g., logits and features) are saved. During the multimodal student model’s training, these pre-computed outputs are loaded from disk per batch in a process analogous to loading the dataset itself. As quantified in Table 10, for a typical 16-sample batch, G2D requires only $\approx$0.5more memory and adds merely $\approx$0.15to the execution time compared to a standard joint-training baseline. Therefore, in resource-constrained environments, if a traditional joint-training model is feasible, G2D is also readily viable by performing the one-time teacher training first and then training the student.
A.7 Additional Ablation Studies
Table 11: G2D vs. Missing Modality Methods on IEMOCAP
Miss Rate Joint-Train CRA [34] MMIN [53] CPM-Net [49] TATE [48] G2D
0% 75.51 76.21 74.94 58.00 69.92 77.19
20% 69.06 67.34 69.36 53.65 63.22 71.49
40% 61.09 57.04 63.30 51.01 60.36 65.10
60% 52.41 43.22 57.52 47.38 57.99 61.50
A.7.1 Learning with Missing Modalities
To evaluate the robustness of G2D with incomplete data, we conduct experiments on the IEMOCAP dataset with randomly missing modalities, generating miss rate masks from 0% to 60% following the setup in MLA [51]. As demonstrated in Table 11, G2D consistently outperforms several state-of-the-art methods designed specifically for this task across all tested rates. This superior performance, even as data becomes highly sparse, suggests that by mitigating modality imbalance and fostering well-rounded representations, our framework learns more resilient features that are less dependent on any single data source, making it inherently more robust when modalities are unavailable.
Table 12: Effect of $\alpha$ and $\beta$ in G2D with SMP
Dataset ($\mathbf{\alpha}$, $\mathbf{\beta}$) weights
(0, 0) (0.25, 0.75) (0.5, 0.5) (0.75, 0.25) (1, 0) (0, 1) (1, 1)
CREMA-D 80.78 84.41 84.95 84.81 82.39 84.68 85.89
UR-Funny 63.58 64.79 64.29 64.29 64.59 64.69 65.49
A.7.2 Effect of $\alpha$ and $\beta$ in G2D.
$\alpha$ and $\beta$ denote the weighting coefficients of feature loss and logit loss, respectively in the proposed G2D loss. Table 12 evaluates the effect of changing the weightage of feature loss and logit loss on G2D. Assigning full weight to both losses ($\alpha=1,\beta=1$) yields the best overall performance, highlighting their combined importance in multimodal learning. In contrast, removing both losses ($\alpha=0,\beta=0$) significantly reduces performance, confirming their necessity. While different weight combinations impact results, incorporating both losses with higher weight leads to greater improvements across datasets.